# Beyond Dialogue: A Profile-Dialogue Alignment Framework Towards General Role-Playing Language Model

### v3

@@ Introduction
Model Character ↑ Style ↑ Emotion ↓ Relationship ↓ Personality ↑ Avg. ↑
w/o Char. 73.75 ± 1.24 81.43 ± 1.40 15.94 ± 0.45 11.75 ± 0.66 67.08 ± 4.56 78.91 ± 1.66
w/o Style. 76.94 ± 1.22 78.99 ± 1.38 15.87 ± 0.45 12.07 ± 0.69 67.25 ± 4.42 79.05 ± 1.63
w/o Emot. 78.69 ± 1.22 80.11 ± 1.38 16.08 ± 0.46 11.82 ± 0.65 67.58 ± 4.92 79.70 ± 1.73
w/o Rela. 77.69 ± 1.29 82.33 ± 1.39 15.81 ± 0.46 12.28 ± 0.68 68.92 ± 4.48 80.17 ± 1.66
w/o Pers. 78.38 ± 1.20 79.34 ± 1.42 16.14 ± 0.45 11.93 ± 0.66 66.17 ± 4.49 79.16 ± 1.64
All 78.67 ± 1.12 82.52 ± 1.33 15.68 ± 0.50 11.22 ± 0.72 69.67 ± 4.27 80.79 ± 1.59
@@
Figure 4 presents a comparison of scores across the five dimensions of the CSERP model, illustrating how the ablation of individual tasks impacts both the overall performance and the specific dimension performances.
@@
Conclusion
This paper introduces a novel general role-playing framework called Beyond Dialogue. We propose a straightforward method that effectively aligns dialogues in specific scenarios with role profiles. This alignment helps to eliminate biases arising from inconsistencies between predefined profiles and the dialogues generated during training. Additionally, we present an innovative prompting mechanism that constructs “Beyond Dialogue” training tasks by generating reasoning processes. It enables fine-grained alignment between role profiles and dialogues at the sentence level. Experimental results demonstrate that our approach enhances the model’s ability to follow predefined profiles across various dimensions of general role-playing, surpassing most general and specialized role-playing baselines. We believe that our framework will inspire further exploration of diverse “Beyond Dialogue” training tasks in the field of general role-playing, thereby enhancing the ability of LLMs to effectively customize proprietary role models.
@@ Appendix A Details of Dataset Construction
Dataset construction process
@@
    Scenario Extract and Chunk Evaluation: Open-source large models are utilized to extract scenarios from the chunks (prompt in Table 10), filtering out those that contain multiple scenarios to avoid discontinuous dialogues. Given the large number of input chunks, using specialized models can be costly. After comparison, we selected Qwen1.5-72B-Chat (Bai, Bai, and et al. 2023) for its balance of performance and efficiency. Additionally, chunks are evaluated to keep only those that reflect the role’s profile (prompt in Table 11), helping to eliminate non-matching chunks and reduce the number of chunks in subsequent inputs.
@@
    Dialogue Extract: Utilizing powerful LLMs to extract dialogues and actions between roles from valid chunks is crucial (prompt in Table 12), as role-playing dialogue data is fundamental for both role-playing training and subsequent derivative tasks. After comparing mainstream models, we selected GPT-4o (gpt-4o-2024-05-13) for its cost-effectiveness. §A details the extraction accuracy and cost of each LLM we evaluated.
@@
    Dialogue Check: To ensure valid role-playing dialogue data, we first focus on forming alternating two-person dialogues, which often leads to scenarios shrinking. This necessitates simultaneous scenario reconstruction and coherence checks to maintain contextual integrity (prompt in Table 13). After ensuring the coherence, we employ GPT-4 to review role portrayal accuracy, discarding any dialogues that conflict with the established profiles (prompt in Table 14).
@@
      Emotion and Relationship: These are aligned using a scale method, with GPT-4o providing a score from 0-10 based on the dialogue content. The Emotion dimension is based on Ekman’s six basic emotions (Ekman 1992): Anger, Disgust, Fear, Happiness, Sadness, and Surprise (prompt in Table 17). The Relationship dimension reflects the intimacy presented in the dialogues within the scenario’s scope (prompt in Table 18).
@@
Comparative Analysis of Dialogue Extraction
This section presents a comparative analysis of LLMs for novel dialogue extraction, focusing on accuracy and cost. The models compared include Qwen1.5-72B-Chat (Bai, Bai, and et al. 2023), ERNIE-Bot (Baidu 2024), deepseek-chat (Deepseek 2024), moonshot-v1-8k (Moonshot 2024), and GPT-4o (OpenAI 2024). Key metrics are dialogue extraction accuracy, omission recall rate, and API costs. Note that Qwen, being open-source, incurs no API costs. A detailed comparison is provided in Table 5.
@@
Statistical Analysis of Beyond Dialogue Role-Playing Dataset
Following the method outlined in §A, we extracted 280 Chinese roles and 31 English roles from 123 Chinese and English novels or scripts. In total, 3,552 sessions of scenario dialogues were obtained, comprising 23,247 dialogue turns, all drawn from authentic dialogues in novels or scripts (Table 6).
@@
Dataset Source Open-source Multi-lingual Multi-turn # Roles # Sessions # Turns # Ave Turns
@@ Appendix B Supplementary Experimental Results
Supplementary Baselines for Main Results
This appendix complements the main results presented in the paper by providing additional insights into the bilingual chatbot baselines used in our evaluation. In addition to our proprietary general baselines, we have also tested newly added open-source general baselines. The proprietary models include GPT-4o, GPT-3.5-Turbo (OpenAI 2024), Moonshot-v1-8k (Moonshot 2024), Yi-Large-Turbo (Wanwu 2024), Deepseek-Chat (Deepseek 2024), Baichuan4 (Baichuan 2024), and Hunyuan (Tencent 2024). The newly tested open-source baselines include Yi-1.5-9B-Chat, GLM-4-9b-chat, Mistral-Nemo-Instruct-2407, and Qwen2-7B-Instruct.
Based on the experimental results, we selected Qwen2-7B-Instruct and Mistral-Nemo-Instruct-2407 for our training due to their balanced performance across both the average score (”Avg.”) and human likeness dimensions. While Yi-1.5-9B-Chat demonstrated a reasonably high average score, its performance in human likeness was notably poor. Its dialogues were excessively rigid and AI-like, which compromised the immersion and authenticity needed in role-playing scenarios.
Moreover, an important factor in selecting Mistral-Nemo-Instruct-2407 was its primary focus on English, in contrast to the other three models—Yi-1.5-9B-Chat, GLM-4-9b-chat, and Qwen2-7B-Instruct—which are predominantly Chinese-language models. Including Mistral in our training set was crucial for ensuring linguistic diversity, thereby enhancing the model’s capability to handle bilingual tasks effectively.
Model Character Style Emotion Relationship Personality Avg. ↑ Human -likeness ↑ Role Choice ↑ Coherence ↑
@@
GPT-4o 74.32 ± 1.15 81.67 ± 1.51 16.31 ± 0.48 12.13 ± 0.66 66.58 ± 4.41 78.83 ± 1.64 67.33 ± 3.95 87.33 ± 3.86 99.67 ± 0.33
GPT-3.5-Turbo 72.26 ± 1.27 73.66 ± 1.73 17.79 ± 0.56 14.17 ± 0.73 66.92 ± 4.85 76.18 ± 1.83 $3$ 4.43 83.00 ± 4.68 97.33 ± 1.17
Moonshot-v1-8k 74.06 ± 1.19 80.64 ± 1.51 16.17 ± 0.47 13.42 ± 0.70 67.00 ± 4.87 78.42 ± 1.75 44.00 ± 4.33 86.67 ± 3.75 99.33 ± 0.46
Yi-Large-Turbo 75.13 ± 1.22 79.18 ± 1.58 16.44 ± 0.49 13.48 ± 0.67 68.25 ± 4.61 78.53 ± 1.72 47.00 ± 4.60 84.33 ± 3.67 92.67 ± 2.39
Deepseek-Chat 75.46 ± 1.14 81.49 ± 1.51 15.92 ± 0.46 12.42 ± 0.63 67.92 ± 4.57 79.30 ± 1.66 52.33 ± 4.95 83.00 ± 4.68 96.67 ± 1.00
Baichuan4 71.82 ± 1.25 76.92 ± 1.52 17.57 ± 0.52 12.30 ± 0.62 67.08 ± 4.75 77.19 ± 1.73 45.33 ± 4.31 82.33 ± 4.49 99.33 ± 0.46
Hunyuan 73.77 ± 1.18 78.75 ± 1.56 17.24 ± 0.48 13.22 ± 0.68 67.00 ± 4.39 77.81 ± 1.66 53.00 ± 4.29 84.33 ± 4.52 98.33 ± 0.84
@@
Yi-1.5-9B-Chat 75.31 ± 1.20 76.78 ± 1.49 16.67 ± 0.52 12.75 ± 0.66 67.42 ± 4.63 78.02 ± 1.70 38.67 ± 4.39 84.00 ± 4.61 92.67 ± 1.79
GLM-4-9b-chat 74.26 ± 1.19 78.40 ± 1.55 17.18 ± 0.50 14.48 ± 0.74 67.17 ± 4.93 77.63 ± 1.78 47.67 ± 4.25 83.33 ± 4.51 99.33 ± 0.46
Mistral-Nemo-Instruct-2407 74.12 ± 1.17 77.04 ± 1.48 17.00 ± 0.43 13.50 ± 0.67 67.00 ± 4.30 77.53 ± 1.61 53.67 ± 4.66 82.67 ± 4.77 74.33 ± 3.77
Qwen2-7B-Instruct 75.39 ± 1.13 77.68 ± 1.65 17.64 ± 0.56 13.43 ± 0.7 67.75 ± 4.44 77.95 ± 1.70 48.00 ± 4.66 83.33 ± 4.48 99.00 ± 0.56
Table 8: Supplementary Main Results of Beyond Dialogue. Bold values indicate the best performance in each group.
Supplementary Result for Ablation Study

### v6

@@ 4.4 Ablation Study
Figure 4 compares scores across the five CSERP dimensions, illustrating how ablation of individual tasks affects both overall and specific performance. Radar charts illustrating these results can be found in § B.3, Figure 10.
Model Character ↑ Style ↑ Emotion ↓ Relationship ↓ Personality ↑ QR
w/o Char. 73.75 ± 1.24 81.43 ± 1.40 15.94 ± 0.45 11.75 ± 0.66 67.08 ± 4.56 45.0 ± 2.87
w/o Style. 76.94 ± 1.22 78.99 ± 1.38 15.87 ± 0.45 12.07 ± 0.69 67.25 ± 4.42 47.67 ± 2.88
w/o Emot. 78.69 ± 1.22 80.11 ± 1.38 16.08 ± 0.46 11.82 ± 0.65 67.58 ± 4.92 48.67 ± 2.89
w/o Rela. 77.69 ± 1.29 82.33 ± 1.39 15.81 ± 0.46 12.28 ± 0.68 68.92 ± 4.48 50.67 ± 2.89
w/o Pers. 78.38 ± 1.20 79.34 ± 1.42 16.14 ± 0.45 11.93 ± 0.66 66.17 ± 4.49 44.67 ± 2.87
All 78.67 ± 1.12 82.52 ± 1.33 15.68 ± 0.50 11.22 ± 0.72 69.67 ± 4.27 56.33 ± 2.86
@@
5 Conclusion
This paper introduces a novel general role-playing framework called Beyond Dialogue. We propose a straightforward method that effectively aligns dialogues in specific scenarios with role profiles. This alignment helps to eliminate distortions arising from biases between predefined profiles and the dialogues generated during training. Additionally, we present an innovative prompting mechanism that constructs “beyond dialogue” training tasks by generating reasoning processes. It enables fine-grained alignment between role profiles and dialogues at the sentence level. In terms of evaluation, we shift from traditional subjective assessments to an objective, efficient, and reproducible method. Experimental results demonstrate that our approach enhances the model’s ability to follow predefined profiles across various dimensions of general role-playing, surpassing most general and specialized role-playing baselines.
Limitations
While our framework addresses the challenge of LLMs adhering to predefined role profiles, real-world scenarios often involve role profiles that change as the dialogue progresses (e.g., Emotions). However, due to the complexity and cost of implementing such tasks, automatically adjusting role profiles based on dialogue evolution and contextual shifts remains an open area for future research.
Additionally, current role-playing models rely on manually predefining scenario-specific profiles. Enabling models to autonomously reason about, align with, and adapt globally-defined role profiles based on evolving conversational contexts — rather than relying on rigid human-crafted definitions — presents another critical direction for advancing dynamic role adaptation.
Furthermore, we recognize that multi-party dialogue introduces additional alignment and evaluation challenges. This scenario requires more sophisticated mechanisms for managing and maintaining role consistency across multiple participants, each with potentially dynamic and evolving profiles. We believe this is an important and underexplored direction for future research.
Ethical Statement
Copyright and Dataset Release. We have ensured that the release of the processed dialogue dataset for academic research, excluding metadata, is in full compliance with current copyright policies. The dataset will be made available under the CC BY-NC 4.0 license, intended solely for non-commercial, academic research purposes.
Use of Human Annotations. Our research involves human annotations for tasks like collecting novel content, character profiles, and validating data accuracy. We employ trained professionals with relevant expertise and ensure fair compensation above the local minimum wage. Informed consent is obtained, and we maintain transparency in the use of annotations. Privacy and ethical standards are prioritized to create a respectful research environment.
Risks. Role-play LLMs trained under the Beyond Dialogue Framework may exhibit only the basic safety alignment of the underlying training model, which means they could potentially generate harmful or toxic content when prompted. As a result, these role-play models are intended solely for research purposes and will require careful alignment for safety in future iterations.
Acknowledgments
This work was sponsored by the National Key Research and Development Program of China (No.2023YFB4606200), Key Program of Science and Technology of Yunnan Province (No.202302AB080020).
@@ Appendix A Details of Dataset Construction
A.1 Dataset construction process
@@
    Scenario Extract and Chunk Evaluation: Open-source large models are utilized to extract scenarios from the chunks (prompt in Table 14), filtering out those that contain multiple scenarios to avoid discontinuous dialogues. Given the large number of input chunks, using specialized models can be costly. After comparison, we selected Qwen1.5-72B-Chat (Yang et al., 2024) for its balance of performance and efficiency. Additionally, chunks are evaluated to keep only those that reflect the role’s profile (prompt in Table 15), helping to eliminate non-matching chunks and reduce the number of chunks in subsequent inputs.
@@
    Dialogue Extract: Utilizing powerful LLMs to extract dialogues and actions between roles from valid chunks is crucial (prompt in Table 16), as role-playing dialogue data is fundamental for both role-playing training and subsequent derivative tasks. After comparing mainstream models, we selected GPT-4o (gpt-4o-2024-05-13) for its cost-effectiveness. § A.2 details the extraction accuracy and cost of each LLM we evaluated.
@@
    Dialogue Check: To ensure valid role-playing dialogue data, we first focus on forming alternating two-person dialogues, which often leads to scenarios shrinking. This necessitates simultaneous scenario reconstruction and coherence checks to maintain contextual integrity (prompt in Table 17). After ensuring the coherence, we employ GPT-4o to review role portrayal accuracy, discarding any dialogues that conflict with the established profiles (prompt in Table 18).
@@
      Emotion and Relationship: These are aligned using a scale method, with GPT-4o providing a score from 0-10 based on the dialogue content. The Emotion dimension is based on Ekman’s six basic emotions (Ekman, 1992): Anger, Disgust, Fear, Happiness, Sadness, and Surprise (prompt in Table 21). The Relationship dimension reflects the intimacy presented in the dialogues within the scenario’s scope (prompt in Table 22).
@@
A.2 Comparative Analysis of Dialogue Extraction
This section presents a comparative analysis of LLMs for novel dialogue extraction, focusing on accuracy and cost. The models compared include Qwen1.5-72B-Chat (Yang et al., 2024), ERNIE-Bot (Sun et al., 2021), deepseek-chat (Bi et al., 2024), moonshot-v1-8k (Team et al., 2025), and GPT-4o (Hurst et al., 2024). Key metrics are dialogue extraction accuracy, omission recall rate, and API costs. Note that Qwen, being open-source, incurs no API costs. A detailed comparison is provided in Table 5.
@@ A.2 Comparative Analysis of Dialogue Extraction
A.3 Statistical Analysis of Beyond Dialogue Role-Playing Dataset
Following the method outlined in § A.1, we extracted 280 Chinese roles and 31 English roles from 123 Chinese and English novels or scripts. In total, 3,552 sessions of scenario dialogues were obtained, comprising 23,247 dialogue turns, all drawn from authentic dialogues in novels or scripts (Table 6).
@@ A.3 Statistical Analysis of Beyond Dialogue Role-Playing Dataset
Dataset Source Open-source Multi-lingual Multi-turn # Roles # Sessions # Turns # Avg Turns
@@
[…]
$\text{UR}=\frac{\sum_{i=1}^{N}\mathbb{I}\left[\exists\,d\in D,\ A_{i}^{(d)}=0\right]}{N}$
The UR represents the proportion of samples (among $N$ total samples) in which at least one dimension $d\in D$ has an alignment status of $A_{i}^{(d)}=0$. Here, $A_{i}^{(d)}\in{0,1}$ indicates whether the $i$-th sample is aligned (1) or unaligned (0) in dimension $d$, and $\mathbb{I}[\cdot]$ is the indicator function that returns 1 if the condition is satisfied.
@@ Appendix B Supplementary Experimental Results
B.1 Subjective vs. Objective Evaluation Results
To assess the impact of evaluation formats on model consistency, we reformulated the evaluation tasks from subjective scoring Wang et al. (2024a); Tu et al. (2024) to objective formats, such as multiple-choice and true/false questions, incorporating Chain-of-Thought (CoT) prompting to minimize variance.
We conducted an experiment with 50 evaluation instances, using mean scores from three human annotators across three settings: subjective consistency scores (Sub), objective questions without CoT (Obj w/o CoT), and objective questions with CoT (Obj), each repeated three times for the five dimensions. We opted to use Krippendorff’s Alpha ($\alpha$) for evaluating the agreement among multiple annotators, including both human evaluators and repeated model annotations.
  1. 1.
    For Sub., since scores are continuous or discrete (e.g., 1–5), we calculate agreement on the final scores.
  2. 2.
    For Obj. and Human, each sample is annotated with multiple classification labels across dimensions(e.g., [1,0,1,0,1]), so we compute $\alpha$ separately for each dimension as a classification task.
Settings Character Style Emotion Relationship Personality
Sub 0.29 0.35 0.53 0.51 0.43
Obj w/o CoT 0.46 0.51 0.65 0.72 0.62
Obj 0.73 0.70 0.80 0.83 0.72
[…]
Table 8: Krippendorff’s Alpha ($\alpha$) values for the evaliation consistency of subjective (Sub.), objective (Obj.), and human (Human) assessments.
As shown in Table 8, the Obj shows significantly higher consistency compared to the Sub (average $\alpha$: 0.76 vs. 0.42), and its consistency is close to that of human annotators. This suggests that combining multiple-choice questions, judgment-based queries, and COT reasoning effectively simulates human evaluative logic, enhancing the stability and reliability of model assessments.
Furthermore, We conducted instance-level comparisons between model-generated scores and human annotations using cosine similarity, computed at the instance level.
Settings Character Style Emotion Relationship Personality
Sub 0.68 0.78 0.85 0.82 0.58
Obj w/o CoT 0.84 0.88 0.80 0.92 0.83
Obj 0.92 0.96 0.94 0.97 0.94
Table 9: Instance-level cosine similarity between model-generated scores and human annotations
Table  9 shows that the Obj achieves the highest average similarity of 0.95 across all CSERP dimensions, closely aligning with human annotations. In contrast, the Sub shows a lower similarity of 0.74, indicating greater deviation from human evaluations. These results demonstrate that the Obj method, with its defined criteria, better aligns with human assessments both at the aggregate and instance levels.
B.2 Cross Model Validation for Main Results
This appendix complements the main results presented in the paper by providing additional insights into the bilingual chatbot baselines used in our evaluation. In addition to our proprietary general baselines, we have also tested newly added open-source general baselines. The proprietary models include Moonshot-v1-8k (Team et al., 2025), Baichuan4 (Yang et al., 2023), and Hunyuan (Sun et al., 2024). The newly tested open-source baselines include Yi-1.5-9B-Chat, GLM-4-9b-chat.
Based on the experimental results, we selected Qwen2-7B-Instruct and Mistral-Nemo-Instruct-2407 for our training due to their strong human-likeness, though their Qualification-Rate lags behind other open-source baselines. We aim for significant improvements in Qualification-Rate with our framework. Moreover, an important factor in selecting Mistral-Nemo-Instruct-2407 was its primary focus on English, in contrast to the other three models — Yi-1.5-9B-Chat, GLM-4-9b-chat, and Qwen2-7B-Instruct — which are predominantly Chinese-language models. Including Mistral in our training set was crucial for ensuring linguistic diversity, thereby enhancing the model’s capability to handle bilingual tasks effectively.
Model Character Style Emotion Relationship Personality Qualification -Rate Human -likeness ↑ Role Choice ↑ Coherence ↑
@@ B.2 Cross Model Validation for Main Results
Moonshot-v1-8k 74.06 ± 1.19 80.64 ± 1.51 16.17 ± 0.47 13.42 ± 0.70 67.00 ± 4.87 45.67 ± 2.88 44.00 ± 4.33 86.67 ± 3.75 99.33 ± 0.46
Baichuan4 71.82 ± 1.25 76.92 ± 1.52 17.57 ± 0.52 12.30 ± 0.62 67.08 ± 4.75 41.33 ± 2.84 45.33 ± 4.31 82.33 ± 4.49 99.33 ± 0.46
Hunyuan 73.77 ± 1.18 78.75 ± 1.56 17.24 ± 0.48 13.22 ± 0.68 67.00 ± 4.39 42.0 ± 2.85 53.00 ± 4.29 84.33 ± 4.52 98.33 ± 0.84
@@
Yi-1.5-9B-Chat 75.31 ± 1.20 76.78 ± 1.49 16.67 ± 0.52 12.75 ± 0.66 67.42 ± 4.63 41.33 ± 2.84 38.67 ± 4.39 84.00 ± 4.61 92.67 ± 1.79
GLM-4-9b-chat 74.26 ± 1.19 78.40 ± 1.55 17.18 ± 0.50 14.48 ± 0.74 67.17 ± 4.93 43.33 ± 2.86 47.67 ± 4.25 83.33 ± 4.51 99.33 ± 0.46
Mistral-Nemo-Instruct-2407 74.12 ± 1.17 77.04 ± 1.48 17.00 ± 0.43 13.50 ± 0.67 67.00 ± 4.30 39.0 ± 2.82 53.67 ± 4.66 82.67 ± 4.77 74.33 ± 3.77
Qwen2-7B-Instruct 75.39 ± 1.13 77.68 ± 1.65 17.64 ± 0.56 13.43 ± 0.7 67.75 ± 4.44 37.33 ± 2.79 48.00 ± 4.66 83.33 ± 4.48 99.00 ± 0.56
Table 10: Main Results of Supplementary Baselines in Beyond Dialogue
In Table 2, we used GPT-4o as the primary evaluation model. To further validate the performance, we also employed the Claude-3-Opus (Anthropic, 2024) model in Table 11, which shares similar capabilities with GPT-4o. While GPT-4o exhibited noticeable declines across several metrics, likely due to its inherent preference for generating its own content, the results from Claude-3-Opus were consistent with those of GPT-4o, thereby confirming the robustness of our evaluation approach.
Model Character Style Emotion Relationship Personality Qualification -Rate Human -likeness ↑ Role Choice ↑ Coherence ↑
Recall ↑ Recall ↑ NMAPE ↓ NMAPE ↓ Precision ↑
General Baselines(Proprietary)
GPT-4o 73.14 ± 1.54 74.76 ± 1.97 15.58 ± 0.57 13.50 ± 0.88 66.00 ± 4.59 41.33 ± 2.84 62.00 ± 8.14 84.00 ± 5.45 99.50 ± 0.50
GPT-3.5-Turbo 70.89 ± 1.65 65.98 ± 2.15 17.16 ± 0.66 16.32 ± 0.95 65.00 ± 5.39 34.67 ± 2.75 30.50 ± 4.32 81.00 ± 6.40 96.00 ± 1.69
Role-play Expertise Baselines
Baichuan-NPC-Turbo 73.38 ± 1.59 75.47 ± 1.66 16.80 ± 0.62 14.58 ± 0.90 64.25 ± 5.45 41.67 ± 2.85 58.00 ± 4.95 83.50 ± 7.08 99.50 ± 0.50
Custom Trained Baselines
Mistral-Nemo-Instruct-2407 74.12 ± 1.17 77.04 ± 1.48 17.00 ± 0.43 13.50 ± 0.67 67.00 ± 4.30 39.0 ± 2.82 53.67 ± 4.66 82.67 ± 4.77 74.33 ± 3.77
+ RP & CC 70.28 ± 1.63 70.49 ± 1.79 16.53 ± 0.56 14.58 ± 0.90 66.75 ± 5.03 42.0 ± 2.85 54.00 ± 4.32 85.50 ± 5.87 83.00 ± 3.17
+ RPA & CC 73.52 ± 1.50 70.70 ± 1.67 15.84 ± 0.56 13.70 ± 0.88 67.88 ± 5.65 45.67 ± 2.88 56.00 ± 4.13 83.00 ± 6.29 91.00 ± 2.40
+ RPA & CC & CSERP 74.58 ± 1.28 78.47 ± 1.45 16.62 ± 0.48 11.38 ± 0.67 69.08 ± 4.46 47.33 ± 2.88 59.00 ± 4.46 87.00 ± 4.73 92.67 ± 1.59
[…]
+ RP & CC 72.56 ± 1.62 72.63 ± 1.68 15.64 ± 0.57 13.90 ± 0.89 65.38 ± 5.03 44.67 ± 2.87 60.00 ± 4.41 80.00 ± 6.77 91.00 ± 1.91
+ RPA & CC 74.33 ± 1.51 75.67 ± 1.71 15.20 ± 0.54 14.68 ± 0.87 67.38 ± 5.18 49.33 ± 2.82 66.00 ± 2.96 86.50 ± 5.25 95.50 ± 1.53
+ RPA & CC & CSERP 76.08 ± 1.45 77.03 ± 1.68 15.29 ± 0.62 12.58 ± 0.93 68.38 ± 4.86 53.0 ± 2.88 65.50 ± 4.13 85.50 ± 5.15 99.00 ± 0.69
Table 11: Supplementary Evaluation Results by Claude-3-Opus (Anthropic, 2024).
B.3 Supplementary Radar Chart of Ablation Results